# arXiv:2603.01665v1[math.DG] 02 Mar 2026

### v1

@@ 1 Introduction
$\mathcal{H}_{\omega}=\left\{u\in\mathcal{C}^{\infty}(X,\mathbb{R})\,|\,\omega_{u}:=\omega+dd^{c}u>0\right\}.$ (1)
plays a central role in Kähler geometry. It is an infinite-dimensional manifold whose tangent space at each point naturally identifies with $C^{\infty}(X)$. At each $u\in\mathcal{H}_{\omega}$, the Mabuchi $L^{2}$-metric [27] on the tangent space at $u$ is
@@
leading to a family of distances $d_{p}$ ($p\geq 1$), and a well defined notion of geodesics between two elements of $\mathcal{H}_{\omega}$. Chen and his collaborators worked intensively in this direction [7, 6, 10, 8, 9] proving in particular that this infinite dimensional space is a path metric space (the metric is the infimum of the length of paths) with $\mathcal{C}^{1,\overline{1}}$ geodesics. Later Chu-Tosatti-Weinkove [11] has established that geodesics are $\mathcal{C}^{1,1}$-regular, which is known to be the optimal regularity thanks to examples of Darvas, Lempert, Vivas [12, 16, 26].
As shown by Chen [7] and [13] the distance $d_{p}$ between two smooth potentials $u_{0},u_{1}$ is realized by the geodesic segment $u_{t}$ :
@@
The understanding of the geometry and topology of $\mathcal{H}_{\omega}$ is crucial for the study of special metrics in Kähler geometry (such as Kähler-Einstein, constant scalar curvature, extremal metrics…). A powerful method to study this infinite-dimensional space is the so called Kähler quantization (or quantization), which is a process to approximate the infinite dimensional space $\mathcal{H}_{\omega}$ and its geometry by finite dimensional negatively curved manifolds. Such a process has been at the core of intense research over the past decades [19, 28, 29, 9, 1, 4, 17].
The "quantization process" requires the Kähler manifold to be projective. More precisely, we consider $(X,\omega)$ a Kähler manifold which carries a holomorphic line bundle $L\rightarrow X$ with a hermitian metric whose curvature is $\omega$. In this setting, $X$ comes with a family of embeddings into projective spaces :
@@
where $H^{0}(X,L^{k})$ is the finite dimensional space of holomorphic sections of the $k$-th tensor powers of $L$. By construction $\frac{1}{k}\Phi_{k}^{*}\omega_{FS}$ lies in the same cohomology class of $\omega$}, hence $\frac{1}{k}\Phi_{k}^{*}\omega_{FS}=\omega+dd^{c}u$, for some $u$ in $\mathcal{H}_{\omega}$. Such smooth function is in the image of the Fubini-Study map
@@
It is well known that a geodesic line in $\hat{\mathcal{H}}(L^{k})$ passing through $H_{0}=PP^{*}$, $P\in GL(N,\mathbb{C})$, writes as $H(t)=P\exp(tA)P^{*}$. We denote by $\lambda_{1}\leq\cdots\leq\lambda_{N_{k}}$ the eigenvalues of $A$. Our next result concerns the globally defined path $\mathbb{R}\ni t\longmapsto\operatorname{FS}_{k}(H(t))$ :
@@
Let $\mathbb{R}\ni t\mapsto H(t)$ be a geodesic in $\hat{\mathcal{H}}(L^{k})$, parametrized with constant length. Assume $A$ is either definite or $\lambda_{1}<0$ and $\lambda_{N_{k}}>0$. Then the path $\mathbb{R}\ni t\longmapsto\operatorname{FS}_{k}(H(t))$ is a quasi-geodesic in $\mathcal{H}_{\omega}$.
@@
Motivated by the study of the space of Kähler potentials on singular varieties [5, 14, 18], we consider the space of potentials with respect to a big form. More precisely, as before we work on $(X,\omega)$ a compact smooth Kähler manifold and we fix another smooth $(1,1)$-form $\theta$, which we assume to be real, closed, semi-positive and big. We say that $\theta$ is big if $\theta$} can be represented by a strictly positive current, or in other words, if there exists $u\in\operatorname{PSH}(X,\theta)$ such that $\theta+dd^{c}u\geq\varepsilon\omega$, for some $\varepsilon>0$. We recall that $\operatorname{PSH}(X,\theta)$ is the set of $\theta$-plurisubharmonic functions, i.e. the set of upper semicontinuous functions $\{u:X\to\mathbb{R}\cup\{-\infty\}$ which locally write as the sum of a smooth and a plurisubharmonic function and such that $\theta_{u}:=\theta+dd^{c}u\geq 0$ in the weak sense of currents.
We set
@@ Acknowledgements
The second author is supported by the ERC grant SiGMA (No. 101125012, PI : Eleonora Di Nezza). She also gratefully acknowledges the University of Melbourne (MATRIX campus) for providing a welcoming and stimulating environment during the research program “Analytic and Geometric Methods on Complex Manifolds" in which this project was completed. Moreover, we thank Sebastian Boucksom, Tamas Darvas and Mattias Jonsson for helpful discussions.
@@ 2 The Fubini-Study map $\operatorname{FS}_{k}$
We mention that Lempert [25, Theorem 1.1] proved that the map $\operatorname{FS}_{k}$ is injective.
@@
The assumption on the eigenvalues of $A$ is optimal as the following example shows. First, we observe that up to work with a large power of $L$ we can suppose that $L$ is very ample and that $H^{0}(X,L)$ has a basis $(s_{1},...,s_{N})$ so that $s_{1},...,s_{p}$ have no common zeros for some $p<N$. Now, given $H(t)=diag(1,...,1,e^{-t},...,e^{-t})$, ie. $\lambda_{1}=\cdots:\lambda_{p}=0<1=\lambda_{p+1}=\cdots=\lambda_{N}$, we can infer that

### v2

# Quantizing Geodesics in Kähler and Sasaki Geometry

@@ 1 Introduction
$\mathcal{H}_{\omega}=\left\{u\in\mathcal{C}^{\infty}(X,\mathbb{R})\,|\,\omega_{u}:=\omega+dd^{c}u>0\right\}$ (1)
plays a central role in Kähler geometry. It is an infinite-dimensional manifold whose tangent space at each point can be identified with $C^{\infty}(X)$. At each $u\in\mathcal{H}_{\omega}$, the Mabuchi $L^{2}$-metric [28] on the tangent space at $u$ is
@@
leading to a family of distances $d_{p}$ ($p\geq 1$), and a well defined notion of geodesics between two elements of $\mathcal{H}_{\omega}$. In fact, $d_{p}$ is defined as the length pseudometric associated to the Finsler $L^{p}$-norm, but one nontrivial point is that it is a genuine metric (i.e. separates points). Chen and his collaborators worked intensively in this direction [7, 6, 10, 8, 9] proving in particular that this infinite dimensional space is a path metric space (the metric is the infimum of the length of paths) with $\mathcal{C}^{1,\overline{1}}$ geodesics. Later Chu-Tosatti-Weinkove [11] has established that geodesics are $\mathcal{C}^{1,1}$-regular, which is known to be the optimal regularity thanks to examples of Darvas, Lempert, Vivas [12, 16, 27]. In particular, geodesics in general leave the space $\mathcal{H}_{\omega}$ ; we will refer to those as weak geodesics.
As shown by Chen [7] and [13] the distance $d_{p}$ between two smooth potentials $u_{0},u_{1}$ is realized by the geodesic $u_{t}$ :
@@
The understanding of the geometry and topology of $\mathcal{H}_{\omega}$ is crucial for the study of special metrics in Kähler geometry (such as Kähler-Einstein, constant scalar curvature, extremal metrics…). A powerful method to study this infinite-dimensional space is the so called Kähler quantization, which is a process to approximate the infinite dimensional space $\mathcal{H}_{\omega}$ and its geometry by finite dimensional negatively curved manifolds. Such a process has been at the core of intense research over the past decades [19, 29, 32, 9, 1, 4, 17].
The "quantization process" requires the Kähler manifold to be projective. More precisely, we consider $(X,\omega)$ a compact Kähler manifold which carries a holomorphic line bundle $L\rightarrow X$ with a Hermitian metric whose curvature is $\omega$. In this setting, $X$ comes with a family of embeddings into projective spaces :
@@
where $H^{0}(X,L^{k})$ is the finite dimensional space of holomorphic sections of the $k$-th tensor power of $L$. By construction $\frac{1}{k}\Phi_{k}^{*}\omega_{FS}$ lies in the same cohomology class as $\omega$}, hence $\frac{1}{k}\Phi_{k}^{*}\omega_{FS}=\omega+dd^{c}u$, for some $u$ in $\mathcal{H}_{\omega}$. Such smooth function is in the image of the Fubini-Study map
@@
It is well known that a geodesic line in $\hat{\mathcal{H}}(L^{k})$ passing through $H_{0}=PP^{*}$, $P\in GL(N,\mathbb{C})$, reads as $H(t)=P\exp(tA)P^{*}$. We denote by $\lambda_{1}\leq\cdots\leq\lambda_{N_{k}}$ the eigenvalues of $A$. Our next result concerns the globally defined path $\mathbb{R}\ni t\longmapsto\operatorname{FS}_{k}(H(t))$ :
@@
Let $\mathbb{R}\ni t\mapsto H(t)$ be a geodesic in $\hat{\mathcal{H}}(L^{k})$, parametrized with constant speed. Assume $A$ is either definite or $\lambda_{1}<0$ and $\lambda_{N_{k}}>0$. Then the path $\mathbb{R}\ni t\longmapsto\operatorname{FS}_{k}(H(t))$ is a quasi-geodesic in $\mathcal{H}_{\omega}$.
@@
It is also worth mentioning that Phong–Sturm [30], Ross–Witt Nyström [31], and Finski [21] (in the more general big setting) show that, given a sequence (in $k$) of geodesic rays on $H^{0}(X,L^{k})$ constructed from carefully chosen norms on $H^{0}(X,L^{k})$ and appropriate speeds, the resulting sequence of quasi-geodesics converges, as $k\rightarrow+\infty$, to a weak geodesic ray. An interesting open question is whether, and under what conditions, the quasi-geodesic $\mathbb{R}\ni t\longmapsto\operatorname{FS}_{k}(H(t))$ (defined on the entire real line) converges to a weak geodesic line.
Motivated by the study of the space of Kähler potentials on singular varieties [5, 14, 18], we consider the space of potentials with respect to a big form. More precisely, as before we work on $(X,\omega)$ a compact smooth Kähler manifold and we fix another smooth $(1,1)$-form $\theta$, which we assume to be real, closed, semi-positive and big. We say that $\theta$ is big if $\theta$} can be represented by a strictly positive current, or in other words, if there exists $u\in\operatorname{PSH}(X,\theta)$ such that $\theta+dd^{c}u\geq\varepsilon\omega$, for some $\varepsilon>0$. When $\theta$ is semi-positive this simply amounts to $\int_{X}\theta^{n}>0$.
We recall that $\operatorname{PSH}(X,\theta)$ is the set of $\theta$-plurisubharmonic functions, i.e. the set of upper semicontinuous functions $\{u:X\to\mathbb{R}\cup\{-\infty\}$ which locally write as the sum of a smooth and a plurisubharmonic function and such that $\theta_{u}:=\theta+dd^{c}u\geq 0$ in the weak sense of currents. We set
@@
We point out that the recent preprint [21, Theorem 1.5 and Proposition 5.2] provides an alternative proof of the convergence stated in Theorem C in the general big (not necessarily semi-positive) setting. As explained in [21, Remark 1.6], the uniform convergence was anticipated. Our arguments yields uniform convergence directly in the semi-positive case.
@@ Acknowledgements
The second author is supported by the ERC grant SiGMA (No. 101125012, PI : Eleonora Di Nezza). She also gratefully acknowledges the University of Melbourne (MATRIX campus) for providing a welcoming and stimulating environment during the research program “Analytic and Geometric Methods on Complex Manifolds" in which this project was completed. Moreover, we thank Sébastien Boucksom, Tamas Darvas, Siarhei Finski and Mattias Jonsson for helpful discussions and comments on a previous draft of the note.
@@ 2 The Fubini-Study map $\operatorname{FS}_{k}$
We mention that Lempert [26, Theorem 1.1] proved that the map $\operatorname{FS}_{k}$ is injective.
@@
The assumption on the eigenvalues of $A$ is optimal as the following example shows. First, we observe that up to work with a large power of $L$ we can suppose that $L$ is very ample and that $H^{0}(X,L)$ has a basis $(s_{1},...,s_{N})$ so that $s_{1},...,s_{p}$ have no common zeros for some $p<N$. Now, given $H(t)=diag(1,...,1,e^{t},...,e^{t})$, ie. $\lambda_{1}=\cdots:\lambda_{p}=0<1=\lambda_{p+1}=\cdots=\lambda_{N}$, we can infer that
@@ 4 Quantization in Sasaki geometry
  • [30] D. H. Phong and J. Sturm. Test configurations for k-stability and geodesic rays. Journal of Symplectic Geometry, 5(2) :221–247, 2007.
  • [31] J. Ross and D. Witt Nyström. Analytic test configurations and geodesic rays. Journal of Symplectic Geometry, 12(1) :125–169, 2014.